# CoDial: Interpretable Task-Oriented Dialogue Systems Through Dialogue Flow Alignment

### v1

@@ Abstract
It is often challenging to teach specialized, unseen tasks to dialogue systems due to the high cost of expert knowledge, training data, and high technical difficulty. To support domain-specific applications—such as law, medicine, or finance—it is essential to build frameworks that enable non-technical experts to define, test, and refine system behaviour with minimal effort. Achieving this requires cross-disciplinary collaboration between developers and domain specialists. In this work, we introduce a novel framework, CoDial (Code for Dialogue), that converts expert knowledge, represented as a novel structured heterogeneous graph, into executable conversation logic. CoDial can be easily implemented in existing guardrailing languages, such as Colang, to enable interpretable, modifiable, and true zero-shot specification of task-oriented dialogue systems. Empirically, CoDial achieves state-of-the-art performance on the STAR dataset for inference-based models and is competitive with similar baselines on the well-known MultiWOZ dataset. We also demonstrate CoDial’s iterative improvement via manual and LLM-aided feedback, making it a practical tool for expert-guided alignment of LLMs in high-stakes domains. [^1] Our code and data are publicly available at https://github.com/radinshayanfar/CoDial.
@@ 1 Introduction
The recent emergence of Large Language Models (LLMs) has transformed the field of Natural Language Processing (NLP), achieving remarkable performance across a wide range of benchmarks. As language models become more widely adopted, experts in high-stakes domains such as law, medicine, and accounting are seeking ways to apply them responsibly to specialized tasks. Many domain experts lack coding skills, so it is important to provide tools that let them define, validate, and refine AI behaviour without writing code Tian et al. (2024); Dahan et al. (2023). Cross-disciplinary collaboration often requires frameworks to possess interpretability and be generalizable to a wide variety of tasks, as domain experts may have limited expertise in programming and often need to understand step-by-step execution of the system. However, creating generalizable conversational systems is challenging due to the complexity of human conversation. Task-Oriented Dialogue (TOD) is an area of research dedicated to accomplishing real-world tasks Qin et al. (2023); Jacqmin et al. (2022). Through these general frameworks, it becomes easier to build intelligent dialogue systems for specific, well-defined tasks across various domains.
Many existing works utilize a schema-oriented framework to enforce complex task structures in TODs (Zhang et al., 2023; Zhao et al., 2023; Mehri and Eskenazi, 2021). These systems convert tasks into a parsable task schema or dialogue flow, allowing experts to specify and integrate new tasks with minimal effort. However, these systems do not satisfy two critical properties: (1) interpretability, the ability to identify how the schema is being used by a language model to arrive at its outputs; and (2) the ability for a domain expert to easily modify the components, which is crucial for many real-world applications. For example, while programmatic representations like those in AnyTOD are technically interpretable, they remain difficult for non-experts to modify, as they lack an intermediary interface that would make the task schema accessible and modifiable. To advance real-world applications and support collaboration with domain experts, we explore a novel approach to create a framework that is easy for an expert to specify, validate, and modify. The contributions of our work are summarized as follows:
  • We propose a novel approach for dialogue flow alignment that aims to minimize human effort requirement through high-level abstraction with strong interpretability. This allows domain experts without programming knowledge to effectively align dialogue systems at inference time.
@@
  • The proposed framework, CoDial, consists of three novel components. The heterogeneous dialogue flow representation allows domain experts to define rich task schemas. The guardrail-grounded code generation pipeline transforms dialogue flows into executable LLM guardrailing program, allowing for flexible control of LLMs in the inference stage. The CoDial human-feedback mechanism incorporates human and LLM feedback to refine and optimize the generated guardrailed conversational models.
  • We demonstrate the effectiveness of our framework on publicly available benchmarks, STAR and MultiWOZ.
    We experiment with different code refinement strategies, by incorporating user feedback through manual and LLM-aided modification.
@@ 2 Related Work
Building generalizable conversational systems is challenging due to the complexity of human conversations, particularly when domain expertise is involved (Chen et al., 2017), leading to a focus on task-oriented systems for specific domains (Jacqmin et al., 2022). While LLMs have demonstrated impressive capability in a wide variety of domains, they struggled with TOD and fell behind if not used properly Hudeček and Dusek (2023). Some research (Zhang et al., 2023; Zhao et al., 2023; Mehri and Eskenazi, 2021) has used a schema-guided approach to generalize TOD systems to unseen tasks. Zhao et al. (2023) viewed the task schema as a program and adopted a neuro-symbolic approach to execute the policy program and control the dialogue flow.
@@
Guardrailing aims to enforce human-imposed constraints to control LLMs in the inference time (Dong et al., 2024; Rebedea et al., 2023; Guardrails AI, ). While it originated from AI safety, they can generally be used to define desired behaviour to constrain. Although traditional dialogue management systems, like Google Dialogflow [^2] https://dialogflow.cloud.google.com, allow rigid modelling of dialogue states, they often lack flexibility to define complex task logic, and it is difficult for a user to further enhance the system. NVIDIA NeMo-Guardrails (Rebedea et al., 2023) is a toolkit that adds programmable guardrails to LLM-based conversational applications without fine-tuning. NeMo-Guardrails employs Colang NVIDIA (2024), a programming language, to establish highly flexible conversational flows and guide LLMs within them. Dong et al. (2024) suggested using neuro-symbolic approaches to guardrail LLMs, where a neural agent (e.g., an LLM) can deal with frequently seen cases, and a symbolic agent can embed human-like cognition through structured knowledge for the rare cases.
@@ 3 Methodology
We introduce CoDial, a novel framework for constructing interpretable dialogue systems without requiring training data or programming expertise, as illustrated in Figure 1. We leverage programmatic LLM guardrailing, which allows flexible control of the behaviour of an LLM in the inference stage, and accordingly, ground TOD to graph-based dialogue flows that define the behaviour. CoDial is composed of three key components: (1) CoDial Heterogeneous Dialogue Flows (CHIEF) that allows domain experts to define the task schema (Section 3.1), (2) Guardrail-Grounded Code Generation (GCG) that automatically generates an executable guardrailing program based on the input dialogue flow (Section 3.2); In this paper, we use Colang NVIDIA (2024) guardrailing language, while any other programmatic guardrailing paradigm can be applied, and (3) CoDial Human Feedback (CHF) that incorporates human/LLM feedback to optimize the generated Colang conversational system (Section 3.3).
@@ 3.1 CoDial Dialogue Flow Representation
We design a structured framework that defines task schema as heterogeneous directed graphs, called CoDial Heterogeneous dIaloguE Flows (CHIEF) representation. Unlike prior work (Mehri and Eskenazi, 2021; Zhang et al., 2023) that define task schema as a homogeneous graph—where the only node type represents user intent, an API return value, or a dialogue state—CHIEF allows for different node or edge types in a heterogeneous manner, supporting more structured task definition. Specifically, CHIEF defines different node types that can define rich metadata and natural language logic to cover a wide range of tasks and domains, inspired by Mosig et al. (2020). To the best of our knowledge, we are the first to frame TOD task schema as a heterogeneous directed graph. We will show that CHIEF representation is compatible with, and can be converted to code for open-source guardrailing libraries. Below, we discuss the main node types and actions in CHIEF.
@@
In addition to nodes, CHIEF supports representing global and fallback actions that are not tied to particular dialogue steps. Global actions can be triggered at any point in the dialogue flow. We also define fallback actions, general responses used when no other action is selected (e.g. Sorry, I can’t help with that).
Algorithm 1 An outline of CoDial’s GCG output.
@@
5:     $h_{2i-1}$leftarrow $\left(h_{2i-2};U_{i}\right)$triangleright User input
@@
9:         $B_{i}$leftarrow IntentResponse(intent)
@@
17:              for each $hv$ in $v^{(s)}_{j}$’s dependents do
18:                  $hv$leftarrow null or false
@@
24:     $B_{i}$, $V^{(H)}$leftarrow NAP(state, $\text{LLM}_{\text{A}}$)
@@
26:         $B_{i}$leftarrow $\text{LLM}_{\text{A}}$($V^{(H)}$)
@@
The defined nodes logically connect with edges. We add a textual condition property to edges to allow conditional branching in dialogue flows. We encode the graphs defined by CHIEF as text in JSON format. The JSON representation consists of a list of nodes and a list of edges (Figure 8). The JSON-encoded CHIEF represented dialogue flow is translated into guardrails code with our automatic code generation pipeline.
3.2 Guardrail-Grounded Code Generation
Guardrailing is a general paradigm to enable inference-stage control over LLMs’ behaviour (Dong et al., 2024; Rebedea et al., 2023). Our work is the first to formulate TOD schema as programmatic guardrailing, which allows highly flexible model behaviour adjustment. Prior work (Zhang et al., 2023) use prompting, which requires prompt engineering and could be inefficient with long prompt contexts. In contrast, guardrailing acts as a proxy between the user and LLM, allowing for black-box LLM alignment and removing the need for model fine-tuning or data collection (Rebedea et al., 2023; Dong et al., 2024).
[…]
Our $\text{prompt}_{\text{GCG}}$ outlines the output $g$, containing $x$’s programmatic logic, as presented in Algorithm 1, with the definitions of the notations provided later in this section. $g$ is a complete and executable guardrailing program powered by an LLM agent $\text{LLM}_{\text{A}}$. The conversation runs within an indefinite while loop, where the agent waits for user input, detects the user’s intent for global actions, predicts the slot variables defined in request nodes (DST component), and selects an action and generates a response (NAP component). In this work, we leverage Colang’s built-in intent detection feature for global actions. Note that DST and NAP are combined and generated by $\text{LLM}_{\text{GCG}}$ as a single program (i.e., $g$). Finally, if the NAP component does not generate a response based on the defined logic (e.g. when user says Goodbye!), the $\text{LLM}_{\text{A}}$ is prompted to choose an action, given the fallback actions and all actions defined in the dialogue flow. Figure 4 illustrates the execution life cycle of the agent.
We denote a conversation between user $U$ and chatbot $B$ as a history of messages, Equation 1, where $U_{i}$ and $B_{i}$ show user’s and chatbot’s $i$-th utterance, respectively. Therefore, the total number of utterances in $h_{2i}$ is $2i$.
@@ 3.2 Guardrail-Grounded Code Generation
Moreover, we define a set of slot variables $V^{(S)}$ that track values for all of the slots defined in all request nodes, and helper variables $V^{(H)}$ that track the state for other (non-request) types of nodes. $V^{(S)}$ and $V^{(H)}$ together form the state of conversation $s=\left(V^{(S)};V^{(H)}\right)$ at each turn, which is used to determine the next action. Please refer to Section A.1 for more details on code generation.
@@
As suggested by Feng et al. (2023), LLM prompting shows promising performance in DST. Therefore, we use a simple prompting approach for DST. We leverage Colang’s Natural Language Description (NLD) feature to extract the value of each slot from the conversation history. For each slot, $\text{prompt}_{\text{GCG}}$ instructs $\text{LLM}_{\text{GCG}}$ to write instructions about extracting the value for that slot, given the history. The NLD instructions then prompt $\text{LLM}_{\text{A}}$ when executed by Colang.
Formally, a slot variable $v^{(s)}_{j}\in V^{(S)}$ is predicted at bot’s turn $i$ as Equation 2, where $p^{(s)}_{j}\in P^{(s)}$ is the prompt generated by $\text{LLM}_{\text{GCG}}$ to extract the value for $v^{(s)}_{j}$.
@@
Since updating a slot may affect the state (e.g. in a search task, modifying the search criteria requires re-executing the search), $\text{LLM}_{\text{GCG}}$ needs to identify the helper variables that need to be invalidated when each slot is updated. We instruct $\text{LLM}_{\text{GCG}}$ to list the helper variables of nodes that are reachable from the updated slot in the graph (i.e., nodes that are direct or indirect children of the slot’s request node). These variables are then reset to null or false, depending on their type, when the slot is updated during execution.
@@
We instruct $\text{LLM}_{\text{GCG}}$ to convert the dialogue flow tree into a conditional logic, consisting of nested if/else statements, to generate a response given the current state. For each node $n_{j}$ in the dialogue flow, an if statement is generated to check whether the conversation is in that node, using $v^{(s)}_{j}$ or $v^{(h)}_{j}$, depending on the type of the node. If the condition holds, the corresponding action for that node is executed; otherwise, the logic proceeds to check its child nodes. For each outgoing edge from a node, the dialogue logic checks whether there is a condition associated with the edge and evaluates whether or not the condition is met. If there is no condition, the edge is followed automatically. Formally, next bot utterance is defined in Equation 3.

### v2

@@ Abstract
Building Task-Oriented Dialogue (TOD) systems that generalize across different tasks remains a challenging problem. Data-driven approaches often struggle to transfer effectively to unseen tasks. While recent schema-based TOD frameworks improve generalization by decoupling task logic from language understanding, their reliance on neural or generative models often obscures how task schemas influence behaviour and hence impair interpretability. In this work, we introduce a novel framework, CoDial (Code for Dialogue), which converts a TOD task schema, represented as a novel structured heterogeneous graph, to programmatic LLM guardrailing code, such as NVIDIA’s Colang, enabling interpretable and efficient alignment of dialogue policies during inference. We introduce two paradigms, $\text{CoDial}_{\text{free}}$ and $\text{CoDial}_{\text{structured}}$, for generating LLM guardrails, and propose a feedback mechanism that integrates human feedback to iteratively improve the generated code. Empirically, CoDial achieves state-of-the-art (SOTA) performance on the widely used STAR dataset and is on par with SOTA on the MultiWOZ dataset, while also providing interpretability. We additionally demonstrate CoDial’s iterative improvement via manual and LLM-aided feedback, making it a practical tool for expert-guided alignment of LLMs in high-stakes domains. [^1] Our code and data are publicly available at https://github.com/radinshayanfar/CoDial.
@@ 1 Introduction
Task-Oriented Dialogue (TOD) systems play a crucial role in a wide range of applications, enabling users to accomplish complex tasks such as flight booking or apartment searching through natural language conversation (Qin et al., 2023). Building TOD systems that are capable of operating across different tasks remains a challenging area of exploration (Jacqmin et al., 2022). Data-driven approaches aim to train models on large corpora of conversations spanning multiple tasks, allowing them to capture task-related conversational patterns. Nonetheless, such models often struggle with generalization: the ability to transfer effectively to new unseen task(s) (Mehri and Eskenazi, 2021).
Many recent TOD works adopt a schema-based approach, decoupling language understanding from task-specific dialogue policy (Zhang et al., 2023; Zhao et al., 2023; Mehri and Eskenazi, 2021). These systems provide an approach to utilize a parsable task schema, often represented as a graph, to encode and enforce complex task logic. Most schema-based methods rely on neural or fully generative-based parsing, which fall short on a key property: interpretability, or the capacity to examine how the schema is utilized by the model to produce specific outputs. In contrast to opaque neural or generative representations, a programmatic formulation allows one to inspect and reason about the decision process, thereby facilitating modification and improvement of the system. Interpretability is also especially crucial in high-stakes domains such as law and medicine, where domain experts with minimal technical knowledge need to specify, validate, and refine AI behaviour (Dahan et al., 2023; Tian et al., 2024). AnyTOD (Zhao et al., 2023) addresses this by treating the task schema as a program to be executed by a language model, thereby offering interpretability. However, this approach requires human knowledge to be defined directly as programming functions, which typically demands greater effort and technical expertise than graph-based representations. This added complexity makes the approach less intuitive and more difficult to adopt, particularly for non-technical users.
To enable zero-shot adaptation of interpretable TOD systems to unseen tasks without requiring direct programming, we propose a novel framework, CoDial (Code for Dialogue). At the core of CoDial, we leverage programmatic Large Language Model (LLM) guardrailing languages, such as Colang (NVIDIA, 2024). We reframe LLM guardrails as the foundation for defining TOD system behaviour. Therefore, CoDial inherits the advantages of programmatic guardrailing, making it interpretable by design and enabling flexible behaviour definition at inference. In this paper, we use Colang NVIDIA (2024), a guardrailing language, to build an interpretable no-code TOD system. Specifically, we convert an input task schema, referred to as a dialogue flow, into Colang code. We introduce two paradigms for generating programmatic guardrails: $\text{CoDial}_{\text{free}}$ and $\text{CoDial}_{\text{structured}}$. The contributions of this work are as follows:
  • We propose a novel approach for effective alignment of dialogue systems to unseen task schemas that is interpretable by design. To our knowledge, we are the first to treat TOD systems as programmatic LLM guardrailing, such as Colang code, and automate its generation. This allows users to align dialogue systems at inference while ensuring interpretability, without requiring manual coding, thereby lowering the barrier for non-technical users and domain experts.
@@
  • The proposed framework, CoDial, consists of three novel components. The heterogeneous dialogue flow representation provides a structure for users to define rich task schemas. The guardrail-grounded code generation pipeline transforms dialogue flows into executable LLM guardrailing programs, allowing for interpretable and flexible control of LLMs in the inference stage. The CoDial human-feedback mechanism incorporates human and LLM feedback to refine and optimize the generated guardrailed conversational models.
  • We demonstrate the effectiveness of our interpretable framework on publicly available TOD benchmarks, STAR and MultiWOZ. Our proposed model achieves new state-of-the-art (SOTA) results on STAR and on par results with SOTA on MultiWOZ in a strict zero-shot setting. We also experiment with different code refinement strategies by incorporating user feedback through manual and LLM-aided modification.
@@ 2 Related Work
While LLMs have demonstrated impressive capability in a wide variety of domains, they struggled with TOD and fell behind if not used properly Hudeček and Dusek (2023). Some research (Zhang et al., 2023; Mehri and Eskenazi, 2021) used a neural schema-guided approach to generalize TOD systems to unseen tasks without interpretability. AnyTOD (Zhao et al., 2023) provided an interpretable neuro-symbolic approach by viewing task schema as a manually-written policy program that controls the dialogue flow. However, beyond the manual coding requirement, AnyTOD also relied on extensive training with highly similar task schemas. As a result, it suffered substantial performance drops when transferred to even slightly different task structures, revealing limited generalization to unseen tasks.
@@
CoDial leverages guardrails to steer LLMs behaviour. Guardrailing aims to enforce human-imposed constraints to control LLMs in the inference time (Dong et al., 2024; Rebedea et al., 2023; Guardrails AI, ). While originating from AI safety, we argue that they can generally be used to define the desired behaviour of LLMs. NVIDIA NeMo-Guardrails (Rebedea et al., 2023) is a toolkit that adds programmable guardrails to LLM-based conversational applications and employs Colang NVIDIA (2024), a programming language, to establish highly flexible conversational flows.
@@
Code generation has made remarkable progress with the introduction of LLMs (Le et al., 2022). Although there are still challenges, such as logical consistency and hallucinations (Liu et al., 2024). LLMs are proficient when in-context examples, documentations, or plans are provided (Jiang et al., 2024). There has been research to improve output by rewriting the input prompt, referred to as prompt optimization (Yuksekgonul et al., 2024). Refer to Section˜A.1 for detailed related work.
@@ 3 Methodology
We introduce CoDial, a novel framework for constructing interpretable TOD systems without requiring training data or manual programming, as illustrated in Figure˜1. A graph-based dialogue flow, defining the behaviour of the TOD system, is the only input of CoDial. The core of our approach is leveraging programmatic LLM guardrailing, which allows interpretable and flexible control over the behaviour of an LLM in the inference stage.
CoDial is composed of three key components: (1) CoDial Heterogeneous Dialogue Flows (CHIEF) that allows users to define the task schema (Section˜3.1), (2) Guardrail-Grounded Code Generation (GCG) that automatically creates a TOD system driven by an executable guardrailing program based on the input dialogue flow (Section˜3.2.2), and (3) CoDial Human Feedback (CHF) that incorporates human/LLM feedback to optimize the generated guardrailing application (Section˜3.3). In this paper, we investigate two code generation paradigms for GCG and use the Colang NVIDIA (2024) guardrailing language, but any other programmatic paradigm can be applied.
@@ 3.1 CoDial Dialogue Flow Representation
We design a structured framework to define rich task schemas, referred to as “dialogue flows”, as heterogeneous directed graphs, called CoDial Heterogeneous dIaloguE Flows (CHIEF) representation. Unlike prior work (Mehri and Eskenazi, 2021; Zhang et al., 2023) that define task schema as an unstructured homogeneous graph—where the single node type represents user intent, an API return value, or a dialogue state—CHIEF allows for different node or edge types in a heterogeneous manner, supporting structured and richer task definition. To the best of our knowledge, we are the first to frame TOD task schema as a heterogeneous directed graph and structure its definition. Specifically, CHIEF provides different node types that can define rich metadata and natural language logic to cover a wide range of tasks and domains, inspired by Mosig et al. (2020). We will show that CHIEF representation can be automatically converted to code for programmatic guardrailing tools. Below, we briefly discuss the main node types and actions in CHIEF. Refer to Section˜A.2 for more details.
@@
The request nodes define variables, hereby referred to as slots, that CoDial tracks throughout the conversation (e.g. the departure location in a taxi booking task). When a conversation reaches this node, the system will request information specified by the slots.
@@
This node specifies a call to an external function within a dialogue flow. External actions enable the designer to execute complex logics.
@@
This node defines a template for providing information to the user (e.g. Your taxi is booked with reference number [ref_no]), and an optional follow-up question (e.g. Do you confirm the booking?).
@@
CHIEF supports global and fallback actions that are not tied to particular dialogue steps. Global actions can be triggered at any point in the dialogue flow (e.g. responding to a greeting). We also define fallback actions, general responses used when no other action is selected (e.g. Sorry, I can’t help with that).
The defined nodes logically connect with edges. We add a textual condition property to edges to allow conditional branching in dialogue flows. We encode the graphs defined by CHIEF as text in JSON format (Figure˜10). The JSON-encoded CHIEF representation is translated into programmatic guardrails with GCG, described below.
3.2 Guardrail-Grounded Code Generation
Guardrailing is a general paradigm to define the flow of conversational systems and enable inference-stage control over LLMs’ behaviour (Dong et al., 2024; Rebedea et al., 2023). Unlike neural models, programming codes are inherently interpretable. Therefore, programmatic guardrailing allows interpretable and flexible behaviour definition in conversational systems. Our work is the first to formulate TOD system as programmatic guardrailing and automate its generation, removing the need and technical barrier of programming, while ensuring interpretability.
[…]
We investigate two different paradigms for implementing $\text{prompt}_{\text{GCG}}$ in GCG. In the first paradigm, denoted as $\text{CoDial}_{\text{free}}$, $\text{prompt}_{\text{GCG}}$ provides LLM with the syntax and semantic rules of the guardrailing language. Because several code implementations may be valid for a given problem, this paradigm leaves $\text{LLM}_{\text{GCG}}$ free to design a guardrailing logic that models the given dialogue flow. The second paradigm directly instructs LLM with a certain dialogue flow modelling approach, specifying the structure of $g$ and how to manage the dialogue, treat each node represented in CHIEF, and implement its equivalent guardrailing code. We denote the latter approach as $\text{CoDial}_{\text{structured}}$. Please refer to Section˜A.2 for more details on code generation.
3.2.1 $\text{CoDial}_{\text{free}}$
Since most LLMs are unfamiliar with guardrailing languages, we include the documentation of our chosen language, Colang, in $\text{prompt}_{\text{GCG}}$. As a preliminary design and due to the large context of the documentation, we hand-pick the most essential chunks to provide $\text{LLM}_{\text{GCG}}$ with a general understanding of Colang’s syntax and semantics.
Figure˜3(a) illustrates an overview of the $\text{prompt}_{\text{GCG}}$ for $\text{CoDial}_{\text{free}}$. The prompt begins with Colang syntax and semantic rules, followed by the input dialogue flow $x$, and concludes with a task description instructing the model to generate Colang code for the flow. The generated code $g$ is an executable guardrailing program that specifies a TOD system aligned to the given CHIEF representation. We also instruct $\text{LLM}_{\text{GCG}}$ to enable Colang’s continuation on unhandled user intent flow to allow $\text{LLM}_{\text{A}}$ to generate output, given fallback actions and all actions defined in the dialogue flow, if the guardrails do not match with the user input in a conversation turn. Figure˜8 shows an example of a generated code in $\text{CoDial}_{\text{free}}$.
Algorithm 1 An outline of $\text{CoDial}_{\text{structured}}$.
@@ 3.2.1 $\text{CoDial}_{\text{free}}$
5:  $h_{2i-1}$ $\leftarrow$ $\left(h_{2i-2};U_{i}\right)$ $\triangleright$ User input
@@
9:   $B_{i}$ $\leftarrow$ IntentResponse(intent)
@@
17:     for each hvhv in $v^{(s)}_{j}$’s dependents do
18:      hvhv $\leftarrow$ null or false
@@
24:  $B_{i}$, $V^{(H)}$ $\leftarrow$ NAP(state, $\text{LLM}_{\text{A}}$)
@@
26:   $B_{i}$ $\leftarrow$ $\text{LLM}_{\text{A}}$($V^{(H)}$)
@@
3.2.2 $\text{CoDial}_{\text{structured}}$
The simple design of $\text{CoDial}_{\text{free}}$ serves as an interpretable baseline where LLMs generate TOD programs from CHIEF representations and language documentation without guidance. Additionally, we propose $\text{CoDial}_{\text{structured}}$, where we explicitly instruct the model on how to structure the code, model the dialogue states, and interpret each CHIEF node type for GCG. Figure˜3(b) shows an overview of $\text{prompt}_{\text{GCG}}$ for $\text{CoDial}_{\text{structured}}$.
Our $\text{prompt}_{\text{GCG}}$ outlines the output guardrailing code $g$, as presented in Algorithm˜1. We define notations later in this section. The conversation runs within an infinite while loop, where the TOD system (1) waits for user input, (2) detects the user’s intent for global actions, (3) predicts the slot variables (Dialogue State Tracking; DST); (4) selects an action and generates a response (Next Action Prediction; NAP). In this work, we leverage Colang’s built-in intent detection feature for global actions. Note that DST and NAP are combined in a single executable program (i.e., $g$). Finally, if the NAP component does not generate a response to the given user utterance (e.g., the conversation strays from the defined logic), $\text{LLM}_{\text{A}}$ is directly prompted to choose from all available actions, including fallbacks, based on the conversation history. Figure˜4 shows the full execution life cycle.
We denote a conversation between user $U$ and chatbot $B$ as a history of messages, Equation 1, where $U_{i}$ and $B_{i}$ show user’s and chatbot’s $i$-th utterance, respectively. Therefore, the total number of utterances in a conversation history $h_{2i}$ is $2i$.
@@ 3.2.2 $\text{CoDial}_{\text{structured}}$
We define a set of slot variables $V^{(S)}$ that track values for all of the slots defined in request nodes, and helper variables $V^{(H)}$ that track the state for other (non-request) types of nodes. The union of $V^{(S)}$ and $V^{(H)}$ forms the state of conversation $s=\left(V^{(S)};V^{(H)}\right)$ at each turn, which is used to determine the next action.
@@
As suggested by Feng et al. (2023), LLM prompting shows promising performance in DST, so we take a similar prompting approach in this work. For each slot, $\text{LLM}_{\text{GCG}}$ creates explicit instructions to extract the value from the entire conversation history. We leverage Colang’s Natural Language Description (NLD) feature to execute these instructions with $\text{LLM}_{\text{A}}$ and save the value to a slot variable. Formally, a slot variable $v^{(s)}_{j}\in V^{(S)}$ is predicted as Equation 2, where $p^{(s)}_{j}\in P^{(s)}$ is the prompt generated by $\text{LLM}_{\text{GCG}}$ to extract the value for $v^{(s)}_{j}$.
@@
We instruct $\text{LLM}_{\text{GCG}}$ to convert the CHIEF graph $x$ into a conditional logic consisting of nested if/else statements, to generate a response given $s_{i}$, the state of the conversation at turn $i$. Each generated if statement corresponds to a node $n_{j}$ in $x$, and aligns $s_{i}$ to the conversation logic outlined by the CHIEF representation. If an if statement holds, this indicates $s_{i}$ is “at” that node and the corresponding action is executed; otherwise, for each outgoing edge at node $n_{j}$, the system checks for traversal. If there is a natural language condition associated with the edge and the condition is met, or if there is no explicit condition, $\text{LLM}_{\text{A}}$ traverses the graph to the associated target node. Formally, next bot utterance is defined in Equation 3.